Two planes are 2500km from an airport it takes one plane flying 50km/h faster 1 hour and 40 minutes Less than the other plane to reach airport find the speed of the two plane


1
Expert's answer
2021-06-09T08:06:05-0400

The time it takes the first plane to reach airport is:


"t_1 = \dfrac{2500km}{50km/h} = 50h"

The time it takes the second plane to reach airport is then (1 hour and 40 minutes or 1.67 hour more):


"t_2 = 50h + 1.67h = 51.67h"

Then the speed of the secon plane is:


"v_2 = \dfrac{2500km}{51.67h} \approx 48.4km/h"

Answer. 48.4 km/h